# On stabilizers in finite permutation groups

### v3

@@ 2.3. Imprimitive groups
We now explain the general strategy to deal with colorings of imprimitive groups. A similar idea is contained in [Ser96, Proof of Theorem 1.2], but we will go a little further. Let $A$ and $B$ be permutation groups on the finite sets $\Omega_{1}$ and $\Omega_{2}$ respectively. Let $G\leqslant B\wr_{|\Omega_{1}|}A$ act imprimitively on $\Omega=\Omega_{1}\times\Omega_{2}$. Let $x,y\geq 2$ be integers. Let $X=(X_{i})_{i=1}^{x}$ be a $x$-coloring of $\Omega_{1}$, namely $\Omega_{1}=\sqcup_{i=1}^{x}X_{i}$. Moreover, suppose we have $x$ inequivalent $y$-colorings of $\Omega_{2}$, say $Y_{1},\ldots,Y_{x}$. In particular, for each $i=1,\ldots,x$, we have that $\Omega_{2}=\sqcup_{j=1}^{y}Y_{i,j}$. The idea is to use the $y$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$. (In fact, we intentionally used the same index $i$ to count the colors in $X$ and the colorings of $\Omega_{2}$). We will construct a new $y$-coloring of $\Omega$, say $Z=(Z_{j})_{j=1}^{y}$, as follows. For each $j=1,\ldots y$, define
@@
It is easy to see that $(Z_{j})_{j=1}^{y}$ is actually a partition of $\Omega$.
@@
The stabilizer of the coloring $Z$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the colorings $Y_{i}$’s in $B$ and the stabilizer of the coloring $X$ in $A$.
@@
Inequivalent $x$-colorings of $\Omega_{1}$ provide inequivalent $y$-colorings of $\Omega$.
@@
Fix $Y_{1},\ldots,Y_{x}$ colorings of $\Omega_{2}$, and let $X,X^{\prime}$ be two colorings of $\Omega_{1}$. Let $Z$ and $Z^{\prime}$ be the colorings of $\Omega=\Omega_{1}\times\Omega_{2}$ obtained from $X$ and $X^{\prime}$ respectively. If $Z$ and $Z^{\prime}$ lies in the same orbit of $B\wr_{|\Omega_{1}|}A$, then the induced colorings of $Z$ and $Z^{\prime}$ on $\Omega_{1}$ need to be in the same orbit of $A$. But these restrictions are precisely $X$ and $X^{\prime}$, and the proof follows. ∎
@@
We work by induction on the cardinality of $\Omega$. Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $Stab_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$. Then $Stab_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work. It follows that we can suppose that $G$ is transitive. Primitive groups can be handled with Lemma 2.2 and Table 1. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis applied to $\Omega_{2}$, we are in one of the following cases:
@@
    Here $\ell_{2,2}(B)=3$. Using the $2$-asymmetric orbits of $B$, and the $3$-asymmetric $3$-colorings of $\Omega_{1}$, arguing as before we can construct $\ell_{3,3}(A)$ good orbits of $G$ (set $x=3$, $y=2$ in the construction above). In fact, note that orbit lengths are multiplied in the imprimitive action, so the length of each orbit will be bounded by $2\cdot 3=6$. Now $\ell_{3,3}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
@@
    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{2,3}(B)=4$. Using the $3$-asymmetric orbits of $B$, and the $2$-asymmetric $4$-colorings of $\Omega_{1}$, as before we can construct $\ell_{4,2}(A)$ good orbits of $G$ (set $x=4$, $y=2$ in the construction above). The length of each orbit will be bounded by $3\cdot 2=6$. Finally, $\ell_{4,2}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
@@
The cardinality of the subset in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$, when $B$ has some orbit of length exceeding $6$, and the degree of $K$ grows. The bound on the derived length follows easily:
@@
Let $\Delta\subseteq\Omega$ be as provided by Theorem 1.1. Then $Stab_{G}(\Delta)$ can be embedded in a direct product of copies of $\mathop{\mathrm{Sym}}(6)$. Every solvable subgroup of $\mathop{\mathrm{Sym}}(6)$ has derived length at most $3$, and the proof follows. ∎
In a previous version of this paper, Corollary 1.2 was obtained directly, without first bounding the length of the orbits. Indeed, the idea of proving a stronger statement by induction and gluing colorings is flexible enough to give structural results directly (see Theorem 2.16 below, for example). The next example provides a transitive solvable group where no subset gives a metabelian stabilizer:
@@
Since $\mathop{\mathrm{Sym}}(3)\wr_{2}\mathop{\mathrm{Sym}}(2)$ is not supersolvable, neither is $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, and the same example shows that supersolvable set-stabilizers need not to exist. This is an improvement on [Glu24, Example 3.4], which deals with nilpotent set-stabilizers.
The next similar construction involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$. We observe that, if a finite group acts on a set (even if not transitively), then the length of each orbit has to divide the order of the group. So, for example, $\mathrm{A\Gamma L}_{1}(8)$ does not have any orbit of length $5$.
@@ 2.4. A result about $3$-colorings
We have to prove that $\ell_{3,2}(G)\geq 5$. We work by induction on $|\Omega|$, and we can suppose that $G$ is transitive. Also, primitive groups can be handled with Lemma 2.7. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis, we have that $\ell_{3,2}(B)\geq 5$. We argue as in Subsection 2.3, with $x=5$ and $y=3$. Using the $2$-asymmetric $3$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the asymmetric $5$-colorings of $\Omega_{1}$, we can construct $\ell_{5,1}(A)$ inequivalent $2$-asymmetric $3$-colorings of $\Omega$. We have $\ell_{5,1}(A)\geq 5$ from Corollary 2.6, which makes the induction smooth and completes the proof. ∎
The correspondent structural result answers [Bab22, Question 7c].
@@
See [Glu24] for another positive result (with two colors) that emphatizes the importance of the prime $2$. In general, five colors are necessary and sufficient to achieve a trivial stabilizer [Ser96, Theorem 1.2].
@@ 2.5. CFSG-free results for nonsolvable groups
The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83]. With the trivial exceptions of the alternating and symmetric groups, the same is actually true in the nonsolvable case, but the known proofs of Cameron, Neumann, Saxl and Seress [CNS84, Ser97] require the Classification of Finite Simple Groups. Actually, Cameron [Cam86] proved a stronger statement concerning the number of regular orbits. We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] is strong enough to provide an easier proof of Cameron’s theorem. [^1] Babai pointed out that a slightly weaker observation, which gives an elementary proof of [CNS84], was made by himself in [Bab22, Theorem 15.1].
@@
Let $G$ be a primitive group of degree $n$ not containing $\mathop{\mathrm{Alt}}(n)$. Then the probability that a random orbit on the power set is a regular orbit tends to $1$ as $n$ grows.
@@
With the notation in [Cam86], it is sufficient to prove that $\varepsilon_{1}(G)\to 0$ when $n$ grows. We mimic Cameron’s proof, which uses the notions of base size and minimal degree in a clever way. At the crucial point, he obtains the desired result provided that $|G|\leq c^{\sqrt{n}}$, where $c$ is a constant (see the Remark at the end of page 308). Then, the Classification of Finite Simple groups is used to bound the order of a primitive group, remaining only with Johnson and Hamming groups, which are handled separately. While Babai’s theorem [Bab81] is not enough, [SW16, Corollary 1.6] is sufficient for this job, having a thresold of $d^{\,n^{1/3}(\log n)^{7/3}}$ where $d$ is a constant. ∎
We stress that Babai’s theorem [Bab81] has several applications, and Theorem 2.15 is a nice example where Sun-Wilmes improvement turns out to be crucial. (See [Ebe24] for another recent example.) We are now ready to prove Theorem 1.3, which answers [Bab22, Question 6a]. Again, we prove a slightly stronger statement.
@@
After Theorem 2.15, it is easy to see that such a constant $C$ exists for all primitive groups. Given $C$, we call good the orbits on the power set with the property in the statement. We can suppose that $G$ is transitive. By the Jordan-Hölder theorem, the property of having all nonalternating composition factors of order at most $C$, is a group property preserved by group extensions. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By induction, $B$ admits two good orbits on $\mathcal{P}(\Omega_{2})$ (equivalently, two good $2$-colorings). Arguing as in Subsection 2.3 (set $x=y=2$), using these good orbits of $B$ as colors for $\Omega_{1}$, and the two good $2$-colorings of $\Omega_{1}$, we can construct two inequivalent good orbits of $G$ on $\mathcal{P}(\Omega)$. The proof is complete. ∎
With more work, and using the Classification of Finite Simple Groups, it should be possible to find an explicit version of Theorem 2.16 (it is enough to find $C$ for all primitive groups).
@@ 3. Primitive groups with solvable stabilizer
Therefore, we will focus on primitive groups, with the goal of proving Theorem 1.5. This proof is more algebraic, but the combinatorial ideas in the previous section will also play a key role. In the case of primitive solvable groups, Theorem 1.5 is [LS24, Theorem 1.1]. This part is really about finding a good centralizer in every irreducible solvable linear group, which was done by Yang [Yan09] using ideas in [MW04]. In general, as explained by Li and Zhang in [LZ11, Sec. 2], it is not hard to see that a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
@@ 3.1. Almost simple groups
In an almost simple group $G$, the socle $T$ is the unique minimal normal subgroup. In fact, let $N\neq T$ be another minimal normal subgroup. Then $[N,T]=1$ and so $N\subseteq C_{G}(T)=1$. This implies that the solvable radical of $G$ is trivial and that every solvable subgroup of $G$ is core-free.
@@
Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag. 4]. We have that the maximal subgroup $H$ is an extension of $H_{0}=H\cap G_{0}$ with
$H/H_{0}\cong G_{0}H/G_{0}\cong G/G_{0},$
which is a section of $G/T$ and so of $\mathop{\mathrm{Out}}(T)$. It follows that
@@
To deal with (iii) uniformly, we use the fact that a solvable linear group has derived length bounded by a function on the rank [Dix68]. This and Lemma 3.2 prove Theorem 1.4 for some absolute constant. The (relatively high) constant $10$ is attained by the Fischer group $Fi_{{}_{23}}$, which makes smooth the remaining computations in (i)-(iii). A more careful inspection of the tables shows that $dl(H_{0})\leq 6$ in the infinite families of (iii), while we use a computer for the remaining single groups. We point out some mistakes in [LZ11, Table 15], and rely instead on the [GAP4] library CTblLib (v1.3.9). [^2] One mistake is in correspondence of $Fi_{{}_{23}}$. In fact, $Fi_{{}_{23}}$ has a unique conjugacy class of solvable maximal subgroups, namely AtlasSubgroup(‘‘Fi23’’,7)), with structure $43^{1+8}.2^{1+6}.3^{1+2}.2S$ 4 (with the ATLAS notation) and derived length $10$. A potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer. With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}.(S4\times 2S4)$. Any group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case. ∎
A weaker version of Theorem 1.4 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple. It would be interesting to find a proof of this curious fact which does not rely on the Classification of Finite Simple Groups.

### v4

@@ 2.3. Imprimitive groups
We now explain the general strategy to deal with colorings of imprimitive groups. Let $A$ and $B$ be permutation groups on the finite sets $\Omega_{1}$ and $\Omega_{2}$ respectively. Let $G\leqslant B\wr_{|\Omega_{1}|}A$ act imprimitively on $\Omega=\Omega_{1}\times\Omega_{2}$. Let $x,y\geq 2$ be integers. Let $X=(X_{i})_{i=1}^{x}$ be a $x$-coloring of $\Omega_{1}$, namely $\Omega_{1}=\sqcup_{i=1}^{x}X_{i}$. Moreover, suppose we have $x$ inequivalent $y$-colorings of $\Omega_{2}$, say $Y_{1},\ldots,Y_{x}$. In particular, for each $i=1,\ldots,x$, we have that $\Omega_{2}=\sqcup_{j=1}^{y}Y_{i,j}$. The point is to use the $y$-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$. (In fact, we intentionally used the same index $i$ to count the colors in $X$ and the colorings of $\Omega_{2}$). We will construct a new $y$-coloring of $\Omega$, say $Z=(Z_{j})_{j=1}^{y}$, as follows. For each $j=1,\ldots y$, define
@@
It is easy to see that $(Z_{j})_{j=1}^{y}$ is a partition of $\Omega$.
@@
The stabilizer of the coloring $Z$ in $G$ can be embedded in the semidirect product between the direct product of the stabilizers of the colorings $Y_{i}$ in $B$ and the stabilizer of the coloring $X$ in $A$. More precisely,
$\cap_{j}\mathop{\mathrm{Stab}_{G}}(Z_{j})\>\leqslant\>\left(\prod_{i}\left(\cap_{j}\mathop{\mathrm{Stab}_{B}}(Y_{i,j})\right)\right)\rtimes\left(\cap_{i}\mathop{\mathrm{Stab}_{A}}(X_{i})\right).$
@@
Fix $Y_{1},\ldots,Y_{x}$ colorings of $\Omega_{2}$, and let $X,X^{\prime}$ be colorings of $\Omega_{1}$. Let $Z$ and $Z^{\prime}$ be the colorings of $\Omega_{1}\times\Omega_{2}$ obtained from $X$ and $X^{\prime}$ respectively. If $X$ and $X^{\prime}$ are inequivalent, then $Z$ and $Z^{\prime}$ are inequivalent.
@@
If $Z$ and $Z^{\prime}$ lies in the same orbit of $B\wr_{|\Omega_{1}|}A$, then the induced colorings of $Z$ and $Z^{\prime}$ on $\Omega_{1}$ need to be in the same orbit of $A$. But these restrictions are precisely $X$ and $X^{\prime}$, and the proof follows. ∎
@@
We work by induction on the cardinality of $\Omega$. Suppose that $\Omega_{1},\ldots,\Omega_{t}$ are the orbits of $G$ on $\Omega,$ and that we can find subsets $\Delta_{j}\subseteq\Omega_{j}$ such that $\mathop{\mathrm{Stab}}_{G}(\Delta_{j})$ has orbits of length at most $6$ on $\Omega_{j}$, for each $j=1,\ldots,t$. Then $\mathop{\mathrm{Stab}}_{G}(\sqcup_{j}\>\Delta_{j})$ will surely work. It follows that we can suppose that $G$ is transitive. The base cases for the induction, primitive groups, can be handled with Lemma 2.2 and Table 1. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis applied to $\Omega_{2}$, we are in one of the following cases:
@@
    Here $\ell_{2,2}(B)=3$. Using the $2$-asymmetric orbits of $B$, and the $3$-asymmetric 33-colorings of $\Omega_{1}$, arguing as before we can construct $\ell_{3,3}(A)$ good orbits of $G$ (set $x=3$, $y=2$ in the construction above). In fact, note that orbit lengths are multiplied in the imprimitive action, so the length of each orbit will be bounded by $2\cdot 3=6$. Now $\ell_{3,3}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
@@
    We can assume $B\cong\mathop{\mathrm{Sym}}(3)$, so that $\ell_{2,3}(B)=4$. Using the $3$-asymmetric orbits of $B$, and the $2$-asymmetric 44-colorings of $\Omega_{1}$, as before we can construct $\ell_{4,2}(A)$ good orbits of $G$ (set $x=4$, $y=2$ in the construction above). The length of each orbit will be bounded by $3\cdot 2=6$. Finally, $\ell_{4,2}(A)\geq\ell_{3,2}(A)\geq 5$ by (2.1) and Lemma 2.7.
@@
The cardinality of the subset in Theorem 1.1 grows linearly with the cardinality of $\Omega$, which is necessary as shown by $B\wr K$, when $B$ has some orbit of length exceeding $6$, and the degree of $K$ grows.
@@
Let $\Delta\subseteq\Omega$ be as provided by Theorem 1.1. Then $\mathop{\mathrm{Stab}}_{G}(\Delta)$ can be embedded in a direct product of copies of $\mathop{\mathrm{Sym}}(6)$. Every solvable subgroup of $\mathop{\mathrm{Sym}}(6)$ has derived length at most $3$, and the proof follows. ∎
The next example provides a transitive solvable group where no subset gives a metabelian stabilizer.
@@
Since $\mathop{\mathrm{Sym}}(3)\wr_{2}\mathop{\mathrm{Sym}}(2)$ is not supersolvable, neither is $H\wr_{2}\mathop{\mathrm{Sym}}(2)$, and the same example shows that supersolvable set-stabilizers need not to exist. This is an improvement on [Glu25, Ex. 3.4], which deals with nilpotent set-stabilizers. The next similar construction involves $\mathrm{A\Gamma L}_{1}(8)$, and provides a transitive solvable group where no subset gives a stabilizer with all orbits of length at most $5$. We observe that, if a finite group acts on a set (even if not transitively), then the length of each orbit has to divide the order of the group. So, for example, $\mathrm{A\Gamma L}_{1}(8)$ does not have any orbit of length $5$.
@@ 2.4. A result about $3$-colorings
We have to prove that $\ell_{3,2}(G)\geq 5$. We work by induction on $|\Omega|$, and we can suppose that $G$ is transitive. Also, primitive groups can be handled with Lemma 2.7. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By the inductive hypothesis, we have that $\ell_{3,2}(B)\geq 5$. We argue as in Subsection 2.3, with $x=5$ and $y=3$. Using the $2$-asymmetric 33-colorings of $\Omega_{2}$ as colors for $\Omega_{1}$, and the asymmetric $5$-colorings of $\Omega_{1}$, we can construct $\ell_{5,1}(A)$ inequivalent $2$-asymmetric 33-colorings of $\Omega$. We have $\ell_{5,1}(A)\geq 5$ from Corollary 2.6, which makes the induction smooth and completes the proof. ∎
The correspondent structural result answers [Bab22, Quest. 7c].
@@
See [Glu25] for another positive result (with two colors) that emphasizes the importance of the prime 22. In general, five colors are necessary and sufficient to achieve a trivial stabilizer [Ser96, Th. 1.2].
@@ 2.5. CFSG-free results for nonsolvable groups
The foundation stone to prove a result like Theorem 1.1 is the fact that a primitive solvable group of large degree has a regular orbit on the power set [Glu83]. With the trivial exceptions of the alternating and symmetric groups, the same is true in the nonsolvable case, but the known proofs of Cameron-Neumann-Saxl and Seress [CNS84, Ser97] require the Classification of Finite Simple Groups. Actually, Cameron [Cam86] proved a stronger statement concerning the number of regular orbits. We observe that a more recent elementary result of Sun and Wilmes [SW15, SW16] is strong enough to provide an easier proof of Cameron’s theorem. [^1] Babai pointed out that a slightly weaker observation, which gives an elementary proof of [CNS84], was made by himself in [Bab22, Th. 15.1].
@@
Let $G$ be a primitive group of degree $n$ not containing $\mathop{\mathrm{Alt}}(n)$. Then the proportion of regular orbits on the power set tends to $1$ as $n\to\infty$.
@@
By the remark near the bottom of page 308 of [Cam86], it suffices to show that $|G|\leq c^{\sqrt{n}}$ for some absolute constant $c$ - this is immediate from [SW16, Cor. 1.6]. ∎
We stress that Babai’s theorem [Bab81] has several applications, and Theorem 2.15 is a nice example where Sun-Wilmes improvement turns out to be crucial. (See [Ebe24] for another recent example.)
We are ready to prove Theorem 1.3. Again, we prove a slightly stronger statement.
@@
After Theorem 2.15, it is easy to see that such a constant $C$ exists for all primitive groups. Given $C$, we call nice the orbits on the power set with the property in the statement. We can suppose that $G$ is transitive. By the Jordan-Hölder theorem, the property of having all nonalternating composition factors of order at most $C$, is a group property preserved by group extensions. Let $\Omega=\Omega_{1}\times\Omega_{2}$, with $|\Omega_{1}|\geq 2$. Let $G\leqslant B\wr_{|\Omega_{1}|}A$, with $A$ primitive on $\Omega_{1}$, and $B$ transitive on $\Omega_{2}$. By induction, $B$ admits two nice orbits on $\mathcal{P}(\Omega_{2})$ (equivalently, two nice $2$-colorings). Arguing as in Subsection 2.3 (set $x=y=2$), using these nice orbits of $B$ as colors for $\Omega_{1}$, and the two nice $2$-colorings of $\Omega_{1}$, we can construct two inequivalent nice orbits of $G$ on $\mathcal{P}(\Omega)$. The proof is complete. ∎
Using the Classification of Finite Simple Groups, it should be possible to find an explicit version of Theorem 2.16 (it is enough to find $C$ for all primitive groups).
@@ 3. Primitive groups with solvable stabilizer
Therefore, we will focus on primitive groups, with the goal of proving Theorem 1.5. This proof is more algebraic, but the combinatorial ideas in the previous section will play a key role. In the case of primitive solvable groups, Theorem 1.5 is [LS25, Th. 1.1]. This part is really about finding a good centralizer in irreducible solvable linear groups, which was done by Yang [Yan09] using ideas in [MW04]. Using the O’Nan-Scott theorem, as explained by Li and Zhang in [LZ11, Sec. 2], it is not hard to see that a primitive nonsolvable group with solvable stabilizer is almost simple or of a specific product type.
@@ 3.1. Almost simple groups
In an almost simple group $G$, the socle $T$ is the unique minimal normal subgroup. This implies that the solvable radical of $G$ is trivial and that every solvable subgroup of $G$ is core-free.
@@
Let $T$ be the socle of $G$, and let $G_{0}$ be the normal subgroup containing $T$ which is described in [LZ11, Pag. 444]. In particular, $G_{0}$ is minimal such that $H_{0}=H\cap G_{0}$ is a maximal subgroup of $G_{0}$. It follows that the maximal subgroup $H$ of $G$ is an extension of $H_{0}$ with
$H/H_{0}\cong G_{0}H/G_{0}=G/G_{0},$
which is a section of $G/T$ and so of $\mathop{\mathrm{Out}}(T)$. Therefore
@@
To deal with (iii) uniformly, we use the fact that a solvable linear group has derived length bounded by a function on the rank [Dix68]. This and Lemma 3.2 prove Theorem 1.4 for some absolute constant. The (relatively high) constant $10$ is attained by the Fischer group $\mathop{\mathrm{Fi}}_{{}_{23}}$, which makes smooth the remaining computations in (i)-(iii). A more careful inspection of the tables shows that $dl(H_{0})\leq 6$ in the infinite families of (iii), while we use a computer for the remaining groups. We point out that there are some mistakes in [LZ11, Tab. 15], and rely instead on the [GAP4] library CTblLib (v1.3.9). [^2] One mistake is in correspondence of $\mathop{\mathrm{Fi}}_{{}_{23}}$. In fact, $\mathop{\mathrm{Fi}}_{{}_{23}}$ has a unique conjugacy class of solvable maximal subgroups, namely AtlasSubgroup(‘‘Fi23’’,7)), with structure $3^{1+8}.2^{1+6}.3^{1+2}.2S_{4}$ (with the ATLAS notation) and derived length $10$. A potential obstruction is given by a certain conjugacy class of solvable maximal subgroups $H$ of the baby monster, which seems to be not available on a computer. With the ATLAS notation [CCN+85], $H$ is isomorphic to $3^{11}].(S_{4}\times 2S_{4})$. Any group of order $3^{11}$ has nilpotency class at most $10$ and derived length at most $4$, showing that $dl(H)\leq 8$ in this case. ∎
A cheap version of Theorem 1.4 says that there exists a constant $C$ such that if $G$ is a finite group with a solvable maximal subgroup $H$, and the derived length of $H$ is more than $C$, then $G$ is not simple. It would be interesting to find an elementary proof of this curious fact.